OpenStudy (anonymous):

Solve for x. -5x + x + 9 = -4x + 12 a.Ø b.{1/7} c.{all reals}

OpenStudy (anonymous):

\[6(x - 2) = 4\]\[(x - 2) = 4/6\]\[(x - 2) = 2/3\]\[x= 2/3+2=2 \frac{2}{3}\]

OpenStudy (yanasidlinskiy):

That does not make sense at all. This is the correct answer: -5x+x+9=-4x+12 ---------> Combine variables on both sides. -4x+9=-4x+12 +4x +4x ---------Both of the x's cancel \[9\neq 12\]